OpenStudy (anonymous):

What is the solution set of 7x^2+3x=0?

OpenStudy (anonymous):

I got {0, -3/7}

OpenStudy (anonymous):

But im not sure if it is right or not?

OpenStudy (solomonzelman):

yes, correct

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 7x^2+3x=0}\) \(\large\color{black}{ \displaystyle 7x^2=-3x}\) \(\large\color{black}{ \displaystyle 7x=-3}\) (x=0 is a solution) \(\large\color{black}{ \displaystyle x=-3/7}\)

OpenStudy (solomonzelman):

so, yes x=0 & x=-3/7